OpenStudy (anonymous):

true or false :When the side of a square is decreased in length by 1 foot, its area is 8 ft2. The length of a side of the square is (equation below)

OpenStudy (zzr0ck3r):

13

OpenStudy (anonymous):

the length of a side of a square is \[1+2\sqrt{2}\]

OpenStudy (lgbasallote):

so what's the question then?

OpenStudy (anonymous):

sorry, i didnt finish it, it's a true or false question

OpenStudy (lgbasallote):

oh lol

OpenStudy (zzr0ck3r):

(1+2sqrt(2)-1)^2 ?= 8

OpenStudy (zzr0ck3r):

(2sqrt(2))^2 = 2^2sqrt(2)^2 = ?

OpenStudy (lgbasallote):

when the side of a square is decreased in length by 1 then the area is 8 so (x-1)^2 = 8 expand the binomial x^2 - 2x + 1 = 8 now subtract 8 from both sides x^2 - 2x + 1 - 8 = 8 - 8 combine like terms x^2 - 2x - 7 = 0 use quadratic formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] substitute the values \[x = \frac{-(-2) \pm \sqrt{(-2)^2 - 4(1)(-7)}}{2(1)}\] do you think that equals \(1 +2\sqrt 2\)?

OpenStudy (zzr0ck3r):

it gave you the size of the square ^^^ this is way to much work for an area of a square to equal 8 we need x^2 = 8 we are told x = (1-2swrt(2) - 1) = x so find x^2

OpenStudy (anonymous):

i'm really confused, can you try to explain step by step?

OpenStudy (zzr0ck3r):

(1+2sqrt(2)-1)^2 = (2sqrt(2))^2 = 2^2*sqrt(2)^2 = 4*2 = 8

OpenStudy (anonymous):

please, is it true or false?

OpenStudy (zzr0ck3r):

what is the area of a square?

OpenStudy (anonymous):

dont the square roots and the squares cancel eachother?

OpenStudy (zzr0ck3r):

yes hence sqrt(2)^2 = 2

OpenStudy (anonymous):

8

OpenStudy (zzr0ck3r):

how do you find the area of the square with side x?

OpenStudy (anonymous):

x^2

OpenStudy (zzr0ck3r):

what is x here?

OpenStudy (anonymous):

1+2squareroot2?

OpenStudy (zzr0ck3r):

but they said minus one from that so now what is it?

OpenStudy (zzr0ck3r):

@willow ?

OpenStudy (anonymous):

2sqrrt2

OpenStudy (anonymous):

so its true?

OpenStudy (zzr0ck3r):

yes

OpenStudy (anonymous):

thanks